*Directions: A triangle has legs of lengths "a" and "b" and hypotenuse of length "c". Find the perimeter. Round to the nearest tenth, if necessary.
Problem: a=5cm, c=13cm

Answer by ozybloke(7)

c2 = a2 + b2
132 = 52 + b2
169 = 25 + b2 [subtract 25 from both sides]
144 = b2
12cm = b
The perimeter is 5 + 12 + 13 = 30cm